OpenStudy (anonymous):

Suppose the wholesale price of a certain brand of medium-sized eggs (p) in dollars per carton is related to the weekly supply (x) in thousands of cartons by the equation 625p^2-x^2 =100 If 25,000 cartons of eggs are available at the beginning of a certain week and the price is falling at the rate of 2c/carton/week, at what rate is the weekly supply falling?

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

So if you set x to equal 25 you can solve for p

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

How old,r u?

OpenStudy (anonymous):

I got p=$1.40

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Solving for x' \[625(2)p-2xx'=0 \] \[\frac{ 1250p }{ 2x}=x'\] \[\frac{ 1250(1.4) }{ 2(0.2)}=5\] So is the weekly supply falling at 5,000 a week?

OpenStudy (anonymous):

Whops I meant to put \[\frac{ 1250(0.2) }{ 2(25) }= 5\]

OpenStudy (xapproachesinfinity):

hmm where did you get 0.2 for p? and why did you substitute x' by 5

OpenStudy (xapproachesinfinity):

2c per carton per week is the rate of change for p meaning p'=2c=0.02$

OpenStudy (xapproachesinfinity):

P changes by unit of cartoon and with weeks x changes only with week

OpenStudy (xapproachesinfinity):

you need to redo your implicit differentiation

OpenStudy (xapproachesinfinity):

you need to rewrite the problem: p changes with how many cartoons we would write is as p(x) x changes by weeks x(w)

OpenStudy (xapproachesinfinity):

the implicit equation is like this \(\huge \rm \frac{d}{dt}\left(x\right) =\frac{625p \frac{d}{dt}\left(p\right)}{x}\)

OpenStudy (xapproachesinfinity):

and you have \(\huge \rm p=\sqrt{\frac{100+x^2}{625}}\)

OpenStudy (xapproachesinfinity):

Check this later @OdinMW

OpenStudy (anonymous):

Hi @xapproachesinfinity thanks for your help! So I changed the equation to \[\frac{ 625p^2 }{ x }=x\] and I should solve for d/dt and not d/dx?

OpenStudy (anonymous):

\[\frac{ (x)1250p \frac{ dx }{ dt } - \frac{ dx }{ dt }(625p^2)}{ x^2 }=\frac{ dx }{ dt }\]

OpenStudy (anonymous):

@freckles hey what do you think?

OpenStudy (xapproachesinfinity):

That won't work! Use the equation i give you

OpenStudy (xapproachesinfinity):

@jim_thompson5910 check this^_^

jimthompson5910 (jim_thompson5910):

This is what I get. x is the supply of eggs in thousands of cartons p is the price in dollars t is the time in weeks "the price is falling at the rate of 2c/carton/week" so dp/dt = -0.02 "If 25,000 cartons of eggs are available at the beginning of a certain week", then we know x = 25. Use this to find the price p \[\Large 625p^2-x^2 =100\] \[\Large 625p^2-25^2 =100\] \[\Large 625p^2-625 =100\] \[\Large 625p^2 =100+625\] \[\Large 625p^2 =725\] \[\Large p^2 = \frac{725}{625}\] \[\Large p^2 = \frac{29}{25}\] \[\Large p = \sqrt{\frac{29}{25}}\] \[\Large p = \frac{\sqrt{29}}{\sqrt{25}}\] \[\Large p = \frac{\sqrt{29}}{5}\] \[\Large p \approx 1.07703\] \[\Large p \approx 1.08\] The price of the carton of eggs is approximately $1.08 ------------------------------------------------------------------------------------ Implicitly differentiate both sides with respect to t Plug in p = 1.08, dp/dt = -0.02 and x = 25 Then isolate dx/dt \[\Large 625p^2-x^2 =100\] \[\Large \frac{d}{dt}[625p^2-x^2] = \frac{d}{dt}[100]\] \[\Large \frac{d}{dt}[625p^2]-\frac{d}{dt}[x^2] = 0\] \[\Large 625*2*p*\frac{dp}{dt}-2x*\frac{dx}{dt} = 0\] \[\Large 625*2*(1.08)*(-0.02)-2(25)*\frac{dx}{dt} \approx 0\] \[\Large -27-50\frac{dx}{dt} \approx 0\] \[\Large -50\frac{dx}{dt} \approx 27\] \[\Large \frac{dx}{dt} \approx \frac{27}{-50}\] \[\Large \frac{dx}{dt} \approx -0.54\] dx/dt is negative to reflect a falling supply

jimthompson5910 (jim_thompson5910):

since x is the number of cartons of eggs in thousands and dx/dt is approximately -0.54, this means the supply shrinks by 0.54*1000 = 540 cartons in this time period

OpenStudy (xapproachesinfinity):

ah yes that's the same thing that i was did I didn't evaluate! i just wanted someone to check the work^_^